OpenStudy (anonymous):

Is anyone good at linear algebra that could help me understand some concepts at a deeper level then just a definition?

OpenStudy (anonymous):

do you want help in groups?

OpenStudy (anonymous):

else what topics in li. algebra?

OpenStudy (anonymous):

vector spaces...........?

OpenStudy (anonymous):

what do you mean in groups? so far we have gotten through vector spaces, subspaces, span, linear independence , and now we are on to linear functions(maps?)

OpenStudy (anonymous):

and i understand vector spaces,subspaces, and linear independence but what exactly is the difference between span and linear independence?

OpenStudy (anonymous):

anyone know?

OpenStudy (anonymous):

Given a set of vectors v_1,...,v_n, a linear combination of those vectors is a vector of the form a_1v_1+...+a_nv_n, for a_1,...,a_n some scalars. The span of the vectors v_1,...,v_n is the set of all the linear combinations of those vectors. I.e., it is the set whose elements are the vectors of the form a_1v_1+...+a_nv_n for all possible values of the coefficients a_1,...,a_n.

OpenStudy (anonymous):

ok so the SPAN tells us the we could get any set of vectors in R^n? where linear indepencence just tells us if there are any simmilar vectors?

OpenStudy (unklerhaukus):

linearly independent vectors of \(\mathbb R^n\) will span a space this is not necessarily all of \(\mathbb R^n\)

OpenStudy (anonymous):

ok so what is the point of span? why is it signifigant?

OpenStudy (anonymous):

i think understand why linear independence is important

OpenStudy (unklerhaukus):

say your in \(\mathbb R^3\), and only two of your vectors are linearly independent, (ie one in a scale multiple of another, ), your span can only be single plane in \(\mathbb R^3\)

OpenStudy (anonymous):

ohhhh ok i see where if 3 were linearly independent the span would be all of R^3? (sorry i dont know how to type the facny R lol)

OpenStudy (anonymous):

\[\mathbb{R} ^{3}\]

OpenStudy (unklerhaukus):

thats right , you could get any where in the solid \ (\mathbb R^3\), by taking a combination of steps in each of the three directions,

OpenStudy (anonymous):

Ah ok i understand! ok so from that we went to Dim's, Bases, and Rank. the only thing im unsure about is what does the basis tell us? i know how to get it and how to get the dimension and rank but not exactly what it means

OpenStudy (unklerhaukus):

if all three vectors were a multiple of each other you could only move along a line , taking combinations of three different sized steps

OpenStudy (anonymous):

thats why you can cut out the other vectors that are linearly dependent right? becasue they are not contributing any new information

OpenStudy (unklerhaukus):

yeah, they will only have scale information,